Can someone help I do not understand problem?
The perimeter of a rectangle is 24m. The length is 3 m more than twice the width. Find the dimensions. What is the length___M?

Answer by checkley79(3341)

P=2L+2W
L=2W+3
42=2(2W+3)+2W
42=4W+6+2W
42=6W+6
6W=42-6
6W=36
W=36/6
W=6 ANS. FOR THE WIDTH.
L=2*6+3
L=12+3=15 ANS. FOR THE LENGTH.
PROOF:
42=2*15+2*6
42=30+12
42=42